OpenStudy (anonymous):

Angle θ lies in the second quadrant, and sin θ = 3/5. cos θ = tan θ =

geerky42 (geerky42):

You mean \(\sin\theta = 3/5\)?

OpenStudy (anonymous):

yes lol sorry

OpenStudy (angelicahicks29):

cos is a -4/5 and tan is a -3/4

OpenStudy (er.mohd.amir):

cos is -4/5 and tan is -4/5 because sin is +ve in second quadrant and other is -ve.